find the measure of each angle of a triangle if the ratio of the measure is 2:5:8
Answer by Alan3354(69443)

find the measure of each angle of a triangle if the ratio of the measure is 2:5:8
------------
2+5+8 = 15
180/15 = 12 degs
--------
2*12 = 24 degs
5*12 = 60 degs
8*12 = 96 degs